OpenStudy (anonymous):

Find the value of x: \[x = \frac{1}{1 + \frac{1}{1+ \frac{1}{1 + ...}}}\]

OpenStudy (bahrom7893):

do we do x = 1/(1+x)?

OpenStudy (anonymous):

\[ x = \frac{1}{1 + x} \] and then ...

OpenStudy (bahrom7893):

so x*(x+1)=1 x^2 +x -1=0

OpenStudy (bahrom7893):

x1 = (-1-sqrt(5))/2 x2 = (-1+sqrt(5))/2

OpenStudy (bahrom7893):

wait it's x2 right?

OpenStudy (anonymous):

yeah it can never be negative

OpenStudy (bahrom7893):

lol yea.. i just noticed haha

OpenStudy (anonymous):

hey this that one question that is troubling me for llife. Can any one explain how to solve these types of question

OpenStudy (anonymous):

refer to this its the same concept: nested fractions and nested radicals http://openstudy.com/study#/updates/4f391d36e4b0fc0c1a0df88b